# Quantum-Corrected Thermodynamics of Conformal Weyl Gravity Black Holes: GUP Effects and Phase Transitions

### v1

@@ V QC Thermodynamics of CWGBHs
Figure 6 presents the heat capacity $C$ as a function of the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, revealing critical thermodynamic behavior that characterizes phase transitions in CWGBHs. The color gradient in this contour plot distinctly separates regions of negative and positive heat capacity, corresponding to thermodynamically unstable and stable phases, respectively. For low values of $\gamma$, a sharp transition from negative to positive $C$ occurs as $r_{h}$ increases, marked by a divergence in the heat capacity that signals a second-order phase transition. This behavior becomes less pronounced as $\gamma$ increases, demonstrating that scale-dependent corrections tend to smooth out the phase structure and can delay or suppress thermodynamic instabilities. The critical region near $r_{h}\approx 1.4$ and $\gamma\lesssim 0.2$ identifies a threshold horizon radius beyond which the BH becomes thermodynamically stable.
@@
Figure 6: Heat capacity $C$ as a function of the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed parameters $\beta=2$ and $k=1$. The plot reveals a critical behavior in BH thermodynamics, where the sign change and divergence in $C$ for low $\gamma$ values indicate second-order phase transitions separating thermodynamically unstable (blue) and stable (red) regions.
VI QC JTE in CWGBHs
@@
The JTE represents a fundamental thermodynamic process characterized by constant enthalpy conditions while allowing temperature variations in response to pressure changes. Within the framework of BH thermodynamics, particularly in the extended phase space formalism that treats the cosmological constant as a thermodynamic pressure, JTE provides a powerful tool for examining isenthalpic evolutionary pathways of gravitational systems [88, 35, 89]. The central diagnostic quantity for this analysis is the Joule-Thomson coefficient, which serves as a critical indicator determining whether a BH undergoes cooling or heating during pressure-driven expansion processes under constant enthalpy conditions.
@@
The heat capacity, which plays an essential role in JTE analysis, has been systematically derived through application of the first law of BH thermodynamics using QC expressions for both entropy and temperature, as demonstrated in Eq.(43) and culminating in the comprehensive result given by Eq.(44). This QC heat capacity captures the subtle modifications introduced by CWG geometry and provides the foundation for understanding thermal response properties in the quantum gravitational regime [90].
@@
The Joule-Thomson coefficient $\mu_{J}$ is rigorously defined as the partial derivative of temperature with respect to pressure under constant enthalpy conditions
@@
where $C$ represents the QC heat capacity derived previously, and $V$ denotes the effective thermodynamic volume associated with the BH system. This formulation explicitly reveals the connection between thermal expansion properties and the Joule-Thomson effect, highlighting the role of heat capacity in mediating the thermodynamic response.
@@
$\mu_{J}\approx\frac{8\pi r^{3}\left(kr^{3}+3\beta^{2}\gamma-2\beta\right)}{3\gamma\beta^{2}\left(-\pi^{2}r^{4}-10\right)+2\beta\left(\pi^{2}r^{4}+10\right)+2kr^{3}\left(\pi^{2}r^{4}-2\right)}$ (48)
This complex expression encapsulates the rich interplay between the various CWG parameters ($k$, $\gamma$, $\beta$) and the horizon geometry, revealing how quantum gravitational corrections modify the classical Joule-Thomson behavior. The explicit forms for the derivatives $\partial T_{H}/\partial r_{h}$ and $\partial P_{C}/\partial r_{h}$ emerge directly from the specific QC thermodynamic relations characteristic of the CWG model, yielding analytical or semi-analytical results depending on the computational complexity of the corrected expressions [91].
@@
Comprehensive numerical analysis reveals the intricate behavior of $\mu_{J}$ as a function of the radius of the horizon $r_{h}$ in different parameter regimes, particularly focusing on variations in the scale-dependent parameter $\gamma$ and the cosmological parameter $k$. The transition between the cooling and heating regimes manifests itself as clear sign changes in $\mu_{J}$, providing an unambiguous identification of the boundaries of the thermodynamic phase. These transitions are not merely mathematical artifacts but represent genuine physical phenomena where the BH thermal response fundamentally changes character, capturing the QC phase structure inherent to CWGBHs under JTE conditions.
Figure 7 presents a comprehensive visualization of the JTE coefficient $\mu_{J}$ as a function of the event horizon radius $r_{h}$ for various values of the scale-dependent parameter $\gamma$, with fixed parameters $\beta=2$ and $k=1$. This plot elegantly illustrates the BH thermal response under isenthalpic expansion conditions, where the sign of $\mu_{J}$ serves as the crucial diagnostic: positive values indicate cooling phases during expansion, while negative values correspond to heating phases. For the baseline case $\gamma=0$, the transition from heating to cooling regimes occurs smoothly without dramatic features. However, as $\gamma$ increases, the curves develop pronounced vertical asymptotes that correspond to divergence points in $\mu_{J}$. These divergences mark critical horizon radii where the BH undergoes Joule-Thomson inversion, representing fundamental changes in the thermal response mechanism.
@@
The systematic shift of these inversion points toward larger $r_{h}$ values with increasing $\gamma$ reveals a profound physical insight: the scale-dependent modifications characteristic of CWG effectively delay the onset of cooling behavior and introduce increasingly complex phase structure into the BH thermodynamics. This behavior demonstrates the remarkable sensitivity of BH thermal properties to quantum-inspired corrections and strongly supports the interpretation that the parameter $\gamma$ functions as a control parameter governing both inversion temperature and associated phase transition characteristics. The rich structure revealed in this analysis emphasizes how CWG provides a natural framework for understanding quantum gravitational effects in BH thermodynamics beyond the limitations of classical GR.
@@
Figure 7: JTE coefficient $\mu_{J}$ versus event horizon radius $r_{h}$ for different values of the scale-dependent parameter $\gamma$, with fixed $\beta=2$ and $k=1$. The divergence points of $\mu_{J}$ indicate Joule-Thomson inversion radii, separating regions of heating ($\mu_{J}<0$) from cooling ($\mu_{J}>0$). As $\gamma$ increases, these inversion points shift to larger $r_{h}$, reflecting the impact of scale-dependent geometry on the thermodynamic response of the BH.
VII Gravitational Redshift in CWGBH Geometry

### v2

@@ 5 QC Thermodynamics of CWGBHs
Figure 8 presents the heat capacity $C$ as a function of the event horizon radius $r_{h}$ for several values of the scale-dependent parameter $\gamma$, revealing critical thermodynamic behavior that characterizes phase transitions in CWGBHs. The line plot employs distinct line styles for $\gamma=\{0,0.05,0.10,0.20\}$, clearly showing regions of negative heat capacity (thermodynamically unstable) and positive heat capacity (thermodynamically stable), separated by divergences where $C\to\pm\infty$. The divergence occurs at the condition $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$ in the denominator of Eq. (45). For low values of $\gamma$, a sharp transition from negative to positive $C$ occurs as $r_{h}$ increases, marked by a divergence in the heat capacity that signals a second-order phase transition. This behavior becomes less pronounced as $\gamma$ increases, demonstrating that scale-dependent corrections tend to smooth out the phase structure and can delay or suppress thermodynamic instabilities. The heat capacity divergence disappears entirely when $\gamma$ exceeds the exact critical threshold
$\gamma_{\rm crit}=\frac{2}{3\beta},$ (46)
which gives $\gamma_{\rm crit}=1/3\approx 0.3333$ for $\beta=2$. Beyond this value, $C$ remains finite and positive for all $r_{h}>0$, indicating that quantum geometric corrections effectively stabilize the BH against thermal fluctuations.
@@
Figure 8: Heat capacity $C$ versus event horizon radius $r_{h}$ for different values of the scale-dependent parameter $\gamma=\{0,0.05,0.10,0.20\}$, with fixed parameters $\beta=2$ and $k=1$. The divergences mark second-order phase transitions where the denominator of Eq. (45) vanishes. For $\gamma<\gamma_{\rm crit}=2/(3\beta)=1/3$, the heat capacity transitions from negative (unstable) to positive (stable) as $r_{h}$ increases past the critical radius $r_{\rm crit}=[(2\beta-3\beta^{2}\gamma)/k]^{1/3}$. For $\gamma>\gamma_{\rm crit}$, no divergence exists and the BH is thermodynamically stable throughout.
It is also worth noting that the incorporation of GUP corrections fundamentally modifies the thermodynamic phase structure through systematic modifications to critical points [78, 12, 65]. The quantum-corrected heat capacity exhibits enhanced complexity:
$C_{\text{GUP}}$ $=C\left(1+\frac{\lambda^{2}l_{p}^{2}}{r_{h}^{2}}\right)^{-1}$
$\quad\times\left(1+\frac{\partial}{\partial r_{h}}\left[\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right]\right),$ (47)
where the second term captures quantum gravitational modifications to thermal response properties through the explicit derivative contribution $\partial/\partial r_{h}[\lambda^{2}l_{p}^{2}/(2r_{h}^{2})]=-\lambda^{2}l_{p}^{2}/r_{h}^{3}$ [118, 102, 111].
The critical radius where $C_{\text{GUP}}$ diverges satisfies
$r_{\text{crit,GUP}}=r_{\text{crit}}\sqrt{1+\frac{\lambda^{2}l_{p}^{2}}{r_{\text{crit}}^{2}}},$ (48)
demonstrating systematic shifts in phase transition boundaries under quantum corrections [46, 94, 39]. We note that for a real positive critical radius to exist, the classical condition $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$ requires $\gamma<2/(3\beta)$. For the parameter choice $\beta=2$, $\gamma=0.05$, and $k=1$, this yields a classical critical radius $r_{\text{crit}}\approx 1.34$.
Figure 9: Critical radius shift under GUP corrections as a function of the quantum parameter $\lambda$, calculated with fixed CWG parameters $\beta=2$, $\gamma=0.05$, $k=1$, yielding a classical critical radius $r_{\text{crit}}\approx 1.34$. The systematic increase in critical radius with increasing $\lambda$ demonstrates quantum stabilization effects that delay the onset of thermodynamic instabilities. The monotonic growth reflects the fundamental role of minimal length scales in modifying phase transition boundaries, with quantum corrections effectively expanding the stable thermodynamic regime through geometric uncertainty effects inherent to the GUP framework.
Figure 10: Quantum-enhanced to classical heat capacity ratio $C_{\text{GUP}}/C_{\text{classical}}$ versus horizon radius $r_{h}$ for GUP parameters $\lambda=\{0.5,1.0,1.5\}$, computed with CWG parameters $\beta=2$, $\gamma=0.05$, and $k=1$. The ratio quantifies the relative magnitude of quantum modifications to thermal response properties, exhibiting systematic deviations that become increasingly pronounced at smaller horizon radii approaching the critical value $r_{\text{crit}}\approx 1.34$. The convergence toward unity at large $r_{h}$ confirms the recovery of classical behavior in macroscopic limits, while the enhanced complexity at small scales demonstrates the emergence of quantum gravitational effects that fundamentally alter the thermodynamic phase structure through minimal length corrections.
Figure 11: Comparison of classical and GUP-corrected heat capacities $C$ versus horizon radius $r_{h}$ for GUP parameters $\lambda=\{0.5,1.0,1.5\}$, with CWG parameters $\beta=2$, $\gamma=0.05$, and $k=1$. The black curve represents the classical heat capacity, while colored curves show the GUP-modified values. The plot demonstrates that GUP corrections systematically reduce the magnitude of the heat capacity, with larger $\lambda$ values producing more pronounced suppression. All curves exhibit positive values in the displayed range $r_{h}>r_{\text{crit}}$, corresponding to the thermodynamically stable phase.
Figure 9 quantitatively demonstrates the systematic boundary shifts predicted by the theoretical framework, revealing monotonic increases in critical radii that reflect quantum stabilization mechanisms. For instance, at $\lambda=1.0$, the critical radius shifts from $r_{\text{crit}}\approx 1.34$ to $r_{\text{crit,GUP}}\approx 1.67$, representing approximately a $25\%$ increase in the phase transition boundary. The comprehensive ratio analysis presented in Figure 10 provides direct quantitative validation of the enhanced complexity introduced by quantum corrections, particularly emphasizing the scale-dependent nature of these modifications. Figure 11 further illustrates the direct comparison between classical and GUP-corrected heat capacities, showing that quantum corrections suppress the thermal response while preserving the qualitative thermodynamic behavior.
These modifications become particularly pronounced in the near-Planckian regime where quantum fluctuations fundamentally alter the stability characteristics of gravitational thermodynamic systems [20, 90]. The computational analysis reveals that GUP effects systematically stabilize the thermal phase structure by delaying critical point onset and introducing quantum coherence effects that are entirely absent in semiclassical treatments, establishing clear demarcation between classical and quantum gravitational domains in the thermodynamic parameter space.
6 QC JT Effect in CWGBHs
@@
The Joule-Thomson (JT) effect represents a fundamental thermodynamic process characterized by constant enthalpy conditions while allowing temperature variations in response to pressure changes. Within the framework of BH thermodynamics, particularly in the extended phase space formalism that treats the cosmological constant as a thermodynamic pressure, the JT effect provides a powerful tool for examining isenthalpic evolutionary pathways of gravitational systems [72, 99, 73]. The central diagnostic quantity for this analysis is the JT coefficient, which serves as a critical indicator determining whether a BH undergoes cooling or heating during pressure-driven expansion processes under constant enthalpy conditions.
@@
The heat capacity, which plays an essential role in JT analysis, has been systematically derived through application of the first law of BH thermodynamics using QC expressions for both entropy and temperature, as demonstrated in Eq. (44) and culminating in the result given by Eq. (45). This QC heat capacity captures the subtle modifications introduced by CWG geometry and provides the foundation for understanding thermal response properties in the quantum gravitational regime [7].
@@
The JT coefficient $\mu_{J}$ is rigorously defined as the partial derivative of temperature with respect to pressure under constant enthalpy conditions [137]
@@
We emphasize that the JT effect is a characteristic of real thermodynamic systems with non-trivial equations of state, in contrast to ideal gases for which $\mu_{J}=0$ identically (since $T(\partial V/\partial T)_{P}=V$). For BHs in the extended phase space, the non-vanishing $\mu_{J}$ reflects genuine thermodynamic structure. During isenthalpic expansion ($dH=0$), the volume increases ($\Delta V>0$) and the pressure decreases ($\Delta P<0$), so the sign of $\mu_{J}=(\partial T/\partial P)_{H}$ determines whether the system cools ($\mu_{J}>0$) or heats ($\mu_{J}<0$). The inversion temperature $T_{\rm inv}$ separates these regimes: for $T>T_{\rm inv}$ the BH heats during expansion, while for $T<T_{\rm inv}$ it cools.
@@
where $C$ represents the QC heat capacity derived previously, and $V$ denotes the effective thermodynamic volume associated with the BH system. This formulation explicitly reveals the connection between thermal expansion properties and the JT effect, highlighting the role of heat capacity in mediating the thermodynamic response.
@@
$\mu_{J}\approx\frac{-4\,r_{h}\left(k\,r_{h}^{3}+3\beta^{2}\gamma-2\beta\right)}{3\left(3\beta^{2}\gamma-2\beta\right)}.$ (52)
This expression follows from the corrected pressure $P=(kr_{h}^{3}+3\beta^{2}\gamma-2\beta)/(8\pi r_{h}^{3})$, whose derivative with respect to $r_{h}$ gives $dP/dr_{h}=-3(3\beta^{2}\gamma-2\beta)/(8\pi r_{h}^{4})$, combined with $dT_{H}/dr_{h}=(kr_{h}^{3}+3\beta^{2}\gamma-2\beta)/(2\pi r_{h}^{3})$. A notable structural feature of the corrected $\mu_{J}$ is that its denominator $3(3\beta^{2}\gamma-2\beta)$ is a constant independent of $r_{h}$, in contrast to the previous expression whose denominator contained $r_{h}$-dependent terms involving $\pi^{2}r_{h}^{4}$. This simplification arises because the corrected pressure $P$ has a cleaner functional form, leading to $dP/dr_{h}$ that factors purely in terms of the CWG parameters without additional horizon-radius coupling. The zeros of $\mu_{J}$ occur when $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$, which determines the JT inversion radius. The explicit forms for the derivatives $\partial T_{H}/\partial r_{h}$ and $\partial P_{C}/\partial r_{h}$ emerge directly from the specific QC thermodynamic relations characteristic of the CWG model [150].
As with the thermodynamic quantities in Section 5, the JT coefficient is presented here as a parametric function of $(r_{h},\gamma)$ with fixed $\beta=2$ and $k=1$, which is standard practice for revealing the functional dependence of transport coefficients on horizon geometry. For specific physical configurations, one should evaluate $\mu_{J}$ at self-consistent horizon radii satisfying $B(r_{h})=0$. Table 3 presents numerical values of $\mu_{J}$ computed at these self-consistent horizons, demonstrating that the BH resides in the heating regime ($\mu_{J}<0$) at its event horizon for all displayed parameter combinations.
$\beta$ $\gamma$ $k$ $r_{h}$ $\mu_{J}$
2.0 0.00 1.0 1.3788 0.6337-0.6337
2.0 0.02 1.0 1.3617 0.5965-0.5965
2.0 0.05 1.0 1.3345 0.5355-0.5355
2.0 0.08 1.0 1.3055 0.4667-0.4667
2.0 0.10 1.0 1.2848 0.4154-0.4154
Table 3: JT coefficient $\mu_{J}$ evaluated at self-consistent horizon radii $r_{h}$ determined from $B(r_{h})=0$, with fixed $\beta=2$ and $k=1$. The negative values indicate that the BH is in the heating regime at its event horizon, with temperature increasing under pressure enhancement. The magnitude $|\mu_{J}|$ decreases with increasing $\gamma$, reflecting the moderating influence of scale-dependent corrections on the thermal response.
The JT inversion occurs when the numerator of Eq. (52) vanishes, i.e., when $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$. For a positive inversion radius to exist, we require $2\beta>3\beta^{2}\gamma$, or equivalently $\gamma<2/(3\beta)$. With $\beta=2$, this yields $\gamma<1/3\approx 0.333$. The inversion radii for small $\gamma$ values are:
$r_{\text{inv}}=\left(\frac{2\beta-3\beta^{2}\gamma}{k}\right)^{1/3},$ (53)
giving $r_{\text{inv}}\approx 1.587$ for $\gamma=0$, $r_{\text{inv}}\approx 1.547$ for $\gamma=0.02$, and $r_{\text{inv}}\approx 1.486$ for $\gamma=0.05$. Since the self-consistent horizon radii (Table 3) are smaller than these inversion radii, the BH at its event horizon always resides in the heating phase, transitioning to the cooling phase only at larger radii beyond the inversion point.
The inversion temperature $T_{\rm inv}$ is obtained by evaluating $T_{H}$ at $r_{h}=r_{\rm inv}$:
$T_{\rm inv}=\frac{-3\beta^{2}\gamma+2\beta}{4\pi\,r_{\rm inv}^{2}}+\frac{\gamma}{4\pi}+\frac{k\,r_{\rm inv}}{2\pi}.$ (54)
For $\beta=2$ and $k=1$, this yields $T_{\rm inv}\approx 0.379$ ($\gamma=0$), $T_{\rm inv}\approx 0.373$ ($\gamma=0.02$), $T_{\rm inv}\approx 0.363$ ($\gamma=0.05$), and $T_{\rm inv}\approx 0.344$ ($\gamma=0.10$). As $\gamma$ increases toward $2/(3\beta)$, $T_{\rm inv}$ decreases, narrowing the cooling window until it vanishes at $\gamma_{\rm crit}=2/(3\beta)$.
@@
Comprehensive numerical analysis reveals the intricate behavior of $\mu_{J}$ as a function of the horizon radius $r_{h}$ in different parameter regimes, particularly focusing on variations in the scale-dependent parameter $\gamma$ and the cosmological parameter $k$. The transition between the cooling and heating regimes manifests itself as clear sign changes in $\mu_{J}$, providing an unambiguous identification of the boundaries of the thermodynamic phase. These transitions are not merely mathematical artifacts but represent genuine physical phenomena where the BH thermal response fundamentally changes character, capturing the QC phase structure inherent to CWGBHs under JT conditions.
@@
Figure 12 presents a visualization of the JT coefficient $\mu_{J}$ as a function of the event horizon radius $r_{h}$ for various values of the scale-dependent parameter $\gamma$, with fixed parameters $\beta=2$ and $k=1$. This plot illustrates the BH thermal response under isenthalpic expansion conditions, where the sign of $\mu_{J}$ serves as the crucial diagnostic: positive values indicate cooling phases during expansion, while negative values correspond to heating phases. For all displayed $\gamma$ values, $\mu_{J}$ starts negative at small $r_{h}$, crosses zero at the inversion radius $r_{\text{inv}}$, and becomes positive for $r_{h}>r_{\text{inv}}$, indicating a transition from heating to cooling regimes. As $\gamma$ increases, the inversion radius shifts to smaller values, and for $\gamma\gtrsim 1/3$ no positive inversion radius exists, meaning the BH remains in the heating regime throughout. The shift of these inversion points toward smaller $r_{h}$ values with increasing $\gamma$ reveals that the scale-dependent modifications characteristic of CWG effectively advance the onset of cooling behavior and introduce increasingly complex phase structure into the BH thermodynamics. This behavior demonstrates the remarkable sensitivity of BH thermal properties to quantum-inspired corrections and strongly supports the interpretation that the parameter $\gamma$ functions as a control parameter governing both inversion temperature and associated phase transition characteristics. The rich structure revealed in this analysis emphasizes how CWG provides a natural framework for understanding quantum gravitational effects in BH thermodynamics beyond the limitations of classical GR.
@@
Figure 12: JT coefficient $\mu_{J}$ versus event horizon radius $r_{h}$ for different values of the scale-dependent parameter $\gamma=\{0,0.02,0.05,0.08\}$, with fixed $\beta=2$ and $k=1$. The transition from negative to positive $\mu_{J}$ marks the JT inversion radius $r_{\text{inv}}$, separating regions of heating ($\mu_{J}<0$) from cooling ($\mu_{J}>0$). As $\gamma$ increases, the inversion radius decreases, reflecting the impact of scale-dependent geometry on the thermodynamic response of the BH. For $r_{h}$ values approaching the self-consistent horizon (see Table 3), the coefficient remains negative, indicating the heating regime.
The two-dimensional landscape of the JT coefficient across the $(r_{h},\gamma)$ parameter space is displayed in Figure 13. The contour plot is restricted to the range $r_{h}\in[0.6,\,3.0]$ to focus on the physically relevant region near the inversion boundary; this restriction is necessary because the corrected $\mu_{J}$ in Eq. (52) possesses a constant denominator $3(3\beta^{2}\gamma-2\beta)$, causing $|\mu_{J}|$ to grow as $\sim r_{h}^{4}$ at large $r_{h}$ and saturate the color scale if the full range is displayed. Within this focused window, $\mu_{J}$ ranges from approximately 1-1 to +2+2, as indicated by the colorbar. The resulting contour morphology differs qualitatively from that produced by the former expression, whose $r_{h}$-dependent denominator suppressed the growth at large radii. In Figure 13, the heating regime ($\mu_{J}<0$, blue) occupies the left portion of the plot at small $r_{h}$, separated from the cooling regime ($\mu_{J}>0$, red) by a white inversion band whose position is governed by the condition $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$. This band shifts to smaller $r_{h}$ as $\gamma$ increases from $0.01$ toward $\gamma_{\rm crit}=1/3$, confirming that larger values of the scale-dependent parameter progressively narrow the heating zone. For $\gamma$ approaching $\gamma_{\rm crit}$, the inversion band moves toward $r_{h}\to 0$, and beyond $\gamma_{\rm crit}$ the entire parameter space would be occupied by the heating phase. The sharp color gradient across the inversion boundary reflects the rapid sign reversal of $\mu_{J}$ and underscores the sensitivity of the isenthalpic thermal response to the CWG parameter $\gamma$.
Figure 13: Filled contour plot of the corrected JT coefficient $\mu_{J}$ as a function of the event horizon radius $r_{h}\in[0.6,\,3.0]$ and the scale-dependent parameter $\gamma$, with fixed $\beta=2$ and $k=1$. The diverging blue-white-red colorscheme with $\mu_{J}$ ranging from approximately 1-1 to +2+2 clearly distinguishes the heating regime ($\mu_{J}<0$, blue) from the cooling regime ($\mu_{J}>0$, red). The white band marks the JT inversion boundary $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$, which shifts to smaller $r_{h}$ with increasing $\gamma$. The constant-denominator structure of the corrected $\mu_{J}$ produces a contour morphology that differs from the original expression, with the cooling region extending over a broader area at $r_{h}>r_{\rm inv}$. The inversion boundary vanishes entirely for $\gamma>\gamma_{\rm crit}=2/(3\beta)=1/3$.
7 Gravitational Redshift in CWGBH Geometry